OpenStudy (anonymous):

Solve: 3(4x - 3) = 51 5 4.5 3.5 3

OpenStudy (goformit100):

Just you have to solve for x.

OpenStudy (anonymous):

are you sure or are you just giving me a random answer because you assume I hate you?

OpenStudy (anonymous):

is it C?

OpenStudy (goformit100):

yes

OpenStudy (goformit100):

any more questions ?

OpenStudy (anonymous):

I hope so, and I really do apologize.

OpenStudy (anonymous):

yes. one more for you.

OpenStudy (anonymous):

do you forgive me?

OpenStudy (goformit100):

post new question area

OpenStudy (anonymous):

that was the question, do you forgive me!

OpenStudy (goformit100):

No, your deeds were really hurting.

OpenStudy (anonymous):

what do you mean?

OpenStudy (goformit100):

I can't forgive you sir.

OpenStudy (anonymous):

why not :/

OpenStudy (anonymous):

@goformit100 you blocked me? lol

OpenStudy (goformit100):

I didn't block you sir

OpenStudy (anonymous):

yeah you did sir.

OpenStudy (goformit100):

No.. Didn't. give me the proof.

OpenStudy (anonymous):

um staph with the weird messages

OpenStudy (anonymous):

no need. I know you did, and that's all that counts.

OpenStudy (anonymous):

The goal is to get x by itself. Whatever you do to one side, you have to do the same on the other or it won't be a true equation. To "move" any aspect of the equation, you have to do the opposite. For example, to solve x - 10 = 3 you have to 'get rid of' the 10, since it's being subtracted you add it to both sides: x - 10 = 3 + 10 +10 ------------------ x - 0 =13 Since the 10's equal 0 on the left side, you can drop the 0 and you are left with x. For your problem: There is more than way to solve this, but since the 3 is already factored out I'll start there: 1. Divide both sides by 3 gives you 4x - 3 = 17 2. Move the 3 to the other side, since it's currently being subtracted you add it to both sides. This gives you 4x = 20 3. Divide both sides by 4 and you get x = 5 The other way to do it is to factor the 3 back in, which means you've got 12x - 9 = 51 1. add 9 to both sides to get rid of it and you have 12x = 60. 2. When you divide both sides by 12, you have x = 5.